OpenStudy (nicole143):

@mathmale Hi! Can you tell me how to find if a relation is a function? And whether it's inverse is one too?

OpenStudy (jdoe0001):

keep in mind that only one-to-one functions have inverse functions \(\bf \begin{array}{llll} \textit{unique "x" for every "y"}\\ \textit{one to one function} \end{array} \begin{array}{llll} x&y\\ \hline\\ a&b\\ c&d\\ e&f\\ g&h\\ i&j \end{array}\qquad \begin{array}{llll} \textit{if "x" has repeats}\\ \textit{is not a function} \end{array} \begin{array}{llll} x&y\\ \hline\\ a&b\\ {\color{red}{ c}}&d\\ d&f\\ g&h\\ {\color{red}{ c}}&j \end{array}\\ \quad \\ \begin{array}{llll} \textit{if "y" has repeats, is NOT }\\ \textit{one to one function} \end{array} \begin{array}{llll} x&y\\ \hline\\ a&{\color{red}{ b}}\\ c&d\\ e&f\\ g&{\color{red}{ b}}\\ i&j \end{array}\)